A sztochasztikus folyamatok gyakorlaton való részvétel

dátum:  II. 7. II. 14. II. 21. II. 28. III. 6. III. 13. III. 20.

diákok:

Baksa Dóra  +  −  −  +  +  +  +
Deme Adrienn  +  −  +  +  +  +  +
File Csenge  −  −  +  +  −  +  +
Ignéczi Andrea  +  +  +  +  +  +  −
Kiss Tibor  +  +  +  +  +  +  +
Koreczki Bence  +  +  +  +  +  +  +
Nagy Ágnes  +  +  +  +  +  +  +
Pinczés József  +  +  −  +  +  +  +
Rácz Zsuzsanna Magdolna  +  +  +  +  +  +  −
Szabó Tímea  +  +  −  +  +  +  +
Szikszai Márton  +  −  +  +  +  −  +
Takács Csaba  +  +  +  +  +  +  +
Ungi Gergely  +  +  +  +  −  +  +
Varha Syuzanna  +  +  +  +  +  +  −
Wessels Tobias  +  +  −  −  −  −  −

dátum:  IV. 3. IV. 10. IV. 17. V. 8. V. 15.

diákok:

Baksa Dóra  +  +  +  −  +
Deme Adrienn  +  +  +  +  −
File Csenge  −  +  +  −  −
Ignéczi Andrea  +  +  +  +  −
Kiss Tibor  +  +  +  +  +
Koreczki Bence  +  −  +  +  +
Nagy Ágnes  +  +  +  −  +
Pinczés József  −  +  +  +  +
Rácz Zsuzsanna Magdolna  +  +  +  +  +
Szabó Tímea  +  +  +  +  +
Szikszai Márton  +  +  +  +  +
Takács Csaba  +  +  +  +  −
Ungi Gergely  +  +  −  +  +
Varha Syuzanna  +  +  +  +  +
Wessels Tobias  −  −  −  −  −